Prove that max( f(n), g(n) ) = Ө ( f(n) + g(n) )


1
Expert's answer
2021-11-02T14:43:24-0400

We know that,


"f(n)\le f(n)+g(n)"


and "g(n)\le f(n)+g(n)"


Now, "max(f(n),g(n))\in O(f(n)+g(n))"


"\Rightarrow f(n)+g(n)\le 2max(f(n),g(n))"

Hence, "max(f(n),g(n))\le2max(f(n),g(n))"

"\Rightarrow max(f(n),g(n)) \in \Omega(f(n)+g(n))"

"\Rightarrow max(f(n),g(n))\in \theta(f(n)+g(n))"

Hence, we can conclude that

"max(f(n),g(n))=\theta(f(n)+g(n))"